# Security of AI Agents

### v2

@@ 3.3.2 Remote vulnerabilities
Uncontrolled AI agents can also be a threat to remote services. Modern LLM-based AI agents can interact with the internet via structured API calling. For example, popular AI agent frameworks like LangChain [langchain] provide pre-defined web-query functionality. If the LLM thinks remote resources are needed, it will generate actions for the agent to query remote hosts provided in the agent’s toolchain. This creates the possibility of making the agent a bot for attacking remote hosts. If there are jailbreak attacks that break the system prompt guard and alignment of the LLM, it can generate dangerous actions telling the agent to repeatedly query the same API resource to scan for vulnerabilities on the API server to use in other attacks. Attackers can also use jailbreak attacks to use agents to scrape data from the remote service provider. Since these agents follow actions generated by LLM, their behavior is distinct from regular social bots on the internet [davis2016botornot, cresci2020decade], leading to insufficient detection and early rejection of these jailbroken AI agent bots.
@@
On the other hand, agent planning that relies on an iterative environment feedback loop can be easily repurposed into a bot for performing DoS attacks. When granted access to local resources, the agent’s action planning affects the availability of the local system. Similarly, if the agent’s planning process requires feedback from the external service provider, it will send requests to the API iteratively to find the ideal action. Since the agents perform actions generated by LLMs on the user’s behalf, they follow the same protocol as human users on the internet, leading to remote vulnerabilities.
4 Defenses
In this section, we propose a few defenses for vulnerabilities identified in section 3. We detail each defense, describing the design and preliminary experiment results to convey the feasibility of each defense technique.
4.1 Sessions
@@
When handling requests from multiple users concurrently, web applications face challenges in maintaining the confidentiality and integrity of each user’s interaction data. In these scenarios, effective session management is one of the best practices. Likewise, AI agent services can adopt a similar approach by using sessions as the protection boundary for requests, where all the requests in the same session may share data and states. Web applications often use distributed session management to ensure the scalability with shared data storage. In a distributed session management scheme, each user session is assigned a unique session ID, and the interaction data is stored in a key/value database (KVDB) where the session ID is the key and the interaction data is the value. AI agents can also use the same approach to establish session connections with users, and store the unique session ID and the question-answer history in a KVDB as its working memory. Since the state of the LLM is defined by the change in its input question as in Equation 1, states also serve as the context for subsequent requests.
However, to successfully make use of sessions as a defense in an AI agent setting, there are still some technical questions to be addressed. First, the way to manage the session connection between each user and the agent needs to be carefully considered. Determining which requests belong to the same session is crucial. The agent designer also needs to consider the time to close a session. When closing a session, the agent needs to transfer its working memory from the KVDB to long-term storage for future use, such as improving its model via fine-tuning. Second, the agent has to embed the session ID into the requests to the AI model. When multiple sessions share the same API key to the foundation model, the agent needs to be able to correlate the session it establishes with the user and the session it establishes with the foundation model. Failure to do so may cause the proposed vulnerabilities to remain unsolved.
Another approach in this direction is to formally model the state of the LLM and AI agents as monad. The state transformer monad [launchbury1995state], as shown in LABEL:lst:state_monad, is the standard solution to enable stateful computations, side effects, and system IO in pure, stateless, effect-free, functional languages like Haskell, Isabelle, Coq, etc. Recall from Equation 1: if we view $Q$ and $A$ as types, we can also write it as a function mapping
$StateLLM:Q\to(A,Q)$ (2)
[…]
@@
4.2 Sandbox
@@
A sandbox restricts the capabilities of the agent program. It enforces the limitation on the program’s access to both local and remote resources. In this section, we describe the application of classic access control provided by sandboxes on agent programs.
Access to local resources
The sandbox restricts the agent’s consumption of local resources such as CPU, memory, and storage. It also limits the agent’s access to a sub-file system. Together with session management, it further isolates the sub-file systems between sessions. To demonstrate the necessity of adding a sandbox to AI agents, we designed BashAgent to interact with the operating system (OS) with bash as its tool, which uses gpt-3.5-turbo [gpt35turbo] to understand user instructions and generate actions. BashAgent has two variants BashAgentplain granted with full accessibility and BashAgentsecure constrained in a docker container. Based on the AgentBench [liu2023agentbench] evaluation framework, we collect and design 95 tasks related to system security to check the harmfulness of unconstrained AI agents. We categorize the tasks into confidentiality, integrity, and availability, and check if the LLM would accept the prompts with malicious intent and generate the attacking actions. We show the results of running BashAgentplain in Table I. We found that our agent BashAgentplain accepts the majority (90/95, about $95\%$) of malicious intents and generates the attacking instructions. Out of the 90 generated attacking commands, 76 of them could be executed successfully in an unprotected environment, making the host system extremely vulnerable in all three security aspects. However, once we apply appropriate sandbox configurations, BashAgentsecure successfully defended against all the LLM-generated attacks. BashAgentplain is based on gpt-3.5-turbo, which was aligned with human values by reinforcement learning from human feedback (RLHF) [ouyang2022training] but still struggles to reject malicious intent in the AI agent use case. Therefore, alignment training will not be enough to secure AI agents, and adding limitations on the access to local resources is necessary for complete security.
TABLE I: Unconstrained AI agents (BashAgentplain) will execute dangerous actions generated by the LLM. #Task is the number of tasks we gathered in this category. #Gen is the number of tasks accepted by the LLM and generates attacking actions. #Exec is the number of LLM-generated commends that are executed successfully and compromise the vulnerabilities.
@@
Access to remote resources
A concurrent study SecGPT [wu2024secgpt] aims to isolate interactions between agent programs and remote resources, such as applications and API servers. Their approach focuses on maintaining the integrity of resources by blocking adversarial system prompts originating from malicious external applications. In contrast, our proposed sandbox environment implements controlled access through mechanisms like whitelists, blacklists, and rate limiting in addition to fundamental interaction isolation. This framework allows resource providers to selectively control the extent of access granted to agent programs, ranging from full permission to complete prohibition or limitations to specific subsets of resources. Consequently, our method enhances security by effectively mitigating unwanted access from AI agents and potential threats posed by adversarial inputs to the agent’s underlying LLM.
4.3 Protecting Models for AI Agents
@@
4.3.1 Sessionless models for AI agents
@@
If the AI agent has no notion of sessions, then either the agent must not fine-tune its LLM on private data or the agent must filter out private or malicious data from the query to the AI model.
@@
The first step is to identify this data, e.g., by pattern matching. The next step is to whitewash them into non-sensitive data; for example, by replacing US social security numbers (SSN) with nine random digits. This leaks no information about the specific SSN but still allows the model to learn from the context around the SSN. AI agent applications often require this harmless version of data to be manipulable. For example, in the web shopping [yao2022webshop] scenario, the website may ask to select a payment method in the form of Visa XXXX showing the last four digits of the credit card number. The agent will extract the last four digits from a whitewashed credit card number which has to be converted back to the original plaintext. In this case, the encryption transformation needs to be structure-preserving and information-preserving to text slicing. We can adapt format-preserving encryption (FPE) [bellare2010ffx] for this purpose. By slicing up the plaintext before applying FPE, we can create a ciphertext that can be sliced and decrypted to obtain the original slice of plaintext.
Definition 4.1 (Format-Preserving Encryption for Text Slicing (FPETS)).
@@
A FPETS is an encryption scheme $E$ such that for all possible private messages $m$ and its indices $i,j$,
$E(m[i\ldots j])=E(m)[i\ldots j],i\leq j.$
FPETS allows language models to read and manipulate private data as ciphertext instead of plaintext, therefore preventing privacy leaks. However, whether encrypting data in a natural language prompt alters its semantics to the LLM or not is unknown. To verify this defense method, we design an evaluation framework that prompts the LLM to operate on encrypted data. All agents created for our evaluations contain an encryptor and decryptor as well as necessary information, such as secret keys and the original plaintext, to perform encryption/decryption. They take a natural language prompt from the user as input and pass the prompt to their internal LLM. The internal LLM is given private data ciphertexts through tools or the prompt. After the LLM processes the user’s request, the agent decrypts any ciphertexts in the LLM’s response before returning the response to the user.
@@
As a proof of concept, we first tested encoded strings before encrypted strings. We generate random strings that include digits and both upper case and lower case letters, and encode them with a simple substitution cipher denoted by $E_{1}$. E1 extends the “rotate-by-13” cipher to operate on the character set mentioned above. Since the “rotate-by-13” cipher is simply a character-by-character rotation, “rotate-by-13,” and therefore E1, supports text slicing on encoded strings. We then ask the agent to return the slice of information we want. The agent responds with the decrypted output for us to validate against the original slice of plaintext. We measure the success rate of this evaluation by
$Succ=\frac{N^{\prime}}{N}$ (3)
where $N$ is the total number of tasks and $N^{\prime}$ is the number of tasks where the agent completed a round trip with no error. Let $D_{1}$ denote the decryption scheme corresponding to $E_{1}$. For confidential data $x$, this process can be formulated as
$x=D_{1}(agent(E_{1}(x))).$
For comparison, we also measure and report the success rate of the agent performing the same tasks with the original randomly-generated string. The results of this initial experiment are recorded in Table II. Since the success rate for slicing ciphertexts was similar to the success rate for slicing plaintext, we proceeded with evaluating our encryption defense.
Figure 9: Sessionless AI agents with HE. Tools in this case need to be supported by the HE scheme, like slicing for RHE and addition or multiplication for FHE.
To demonstrate that definition 4.1 works, we implement an SSN agent that manipulates nine-digit SSNs. In our design, all processes involving the plaintext of the SSNs are done outside of the underlying LLM. This ensures that the LLM is never exposed to raw sensitive information. We first provide the agent with an array of four secret keys and map each secret key to its respective randomly-generated SSN. The agent uses this information to encrypt each SSN and store the ciphertexts in an array. We then prompt the agent to return certain groups of an SSN, such as the first three or last four digits. Throughout its reasoning process, the underlying LLM of the agent has no access to the original plaintext of the SSNs. The LLM could at most call a tool to retrieve an SSN’s ciphertext for reasoning based on a fictitious user ID for indexing the array. After the LLM responds with the ciphertext representation of the slice we asked for, the agent replaces the ciphertext within the response with its decrypted value outside of the LLM before returning it to the user. We verified the results by comparing the agent’s response to the actual slice of plaintext we expected. The experimental results are shown in Table II.
@@
Despite an unimpressive success rate, the results showed that gpt-3.5-turbo and gpt-4-turbo were able to understand and respond to queries involving the manipulation of encoded strings. Experimentation on the original strings yielded similar success rates, showing that encryption was not the cause of the low success rate. This means that encrypted data in the prompt does not alter the semantics of the input query.
@@
TABLE II: Results for AI agent with encrypted data. Each agent is evaluated on 100 randomly-generated tasks. “SuccCiph” is the success rate of agent completing the tasks with encrypted data, as in Equation 3. “SuccPlain” is the success rate of the agent completing the same tasks without encrypting the data.
@@
Slicing gpt-3.5-turbo 49.0% 47.0%
Slicing gpt-4-turbo 55.0% 57.0%

### v3

@@ III-C2 Remote vulnerabilities
Uncontrolled AI agents can also be a threat to remote services. Modern LLM-based AI agents can interact with the internet via structured API calling. For example, popular AI agent frameworks like LangChain provide pre-defined web-query functionality. If the LLM thinks remote resources are needed, it will generate actions for the agent to query remote hosts provided in the agent’s toolchain. This creates the possibility of making the agent a bot for attacking remote hosts. If there are jailbreak attacks that break the system prompt guard and alignment of the LLM, it can generate dangerous actions telling the agent to repeatedly query the same API resource to scan for vulnerabilities on the API server to use in other attacks. Attackers can also use jailbreak attacks to use agents to scrape data from the remote service provider. Since these agents follow actions generated by LLM, their behavior is distinct from regular social bots on the internet [davis2016botornot], leading to insufficient detection and early rejection of these jailbroken AI agent bots.
@@
Furthermore, agent planning that relies on an iterative environment feedback can be easily repurposed into a bot for performing DoS attacks. When granted access to local resources, the agent’s action planning affects the availability of the local system. Similarly, if the agent’s planning process requires feedback from the external service provider, it will send requests to the API iteratively to find the ideal action. Since the agents perform actions generated by LLMs on the user’s behalf, they follow the same protocol as human users on the internet, leading to remote vulnerabilities.
IV Defenses
We propose defenses for the vulnerabilities in section III. We describe their design and evaluate their feasibility through experiments and empirical analysis.
IV-A Sessions
@@
When handling requests from multiple users concurrently, web applications face challenges in maintaining the confidentiality and integrity of each user’s interaction data. In these scenarios, effective session management is one of the best practices. Likewise, AI agent services can adopt a similar approach by using sessions as the protection boundary for requests, where all the requests in the same session may share data and states. Web applications often use distributed session management to ensure the scalability with shared data storage. In a distributed session management scheme, each user session is assigned a unique session ID, and the interaction data is stored in a key/value database (KVDB) where the session ID is the key and the interaction data is the value as shown in Figure 6. AI agents can also use the same approach to establish session connections with users, and store the unique session ID and the question-answer history in a KVDB as its working memory. Since the state of the LLM is defined by the change in its input question as in Equation 1, states also serve as the context for subsequent requests.
However, to successfully use sessions as defense in AI agents, technical challenges remain. First, the way to manage the session connection between each user and the agent needs to be carefully considered. Determining which requests belong to the same session is crucial. The agent designer also needs to consider the time to close a session. When closing a session, the agent needs to transfer its working memory from the KVDB to long-term storage for future use, such as improving its model via fine-tuning. Second, the agent has to embed the session ID into the requests to the AI model. When multiple sessions share the same API key to the foundation model, the agent needs to be able to correlate the session it establishes with the user and the session it establishes with the foundation model. Otherwise, the described vulnerabilities will remain.
@@
Another approach in this direction is to formally model the state of the LLM and AI agents as monad. The state transformer monad [launchbury1995state] is the standard solution to enable stateful computations, side effects, and system IO in pure, stateless, effect-free, functional languages like Haskell, Isabelle, Coq, etc. Recall from Equation 1: if we view $Q$ and $A$ as types, we can also write it as a function mapping $StateLLM:Q\to(A,Q)$, which transforms the LLM from an initial state to the next state. Then the formal definition of the state transformer [launchbury1995state] is a parametric form of this function as shown in 1. Since monads are composable [jones1993composing], the state monad is particularly ideal for representing AI agent behaviors such as reasoning and planning. We show a few examples in Figure 7 to demonstrate this idea as an analogy to [launchbury1995state]. We believe future research can build on this framework to derive a formal definition of the state of AI agents. The state monad is defined in a formal type system with type inference that is both sound and complete [schlesinger2012verification], which may facilitate the verification of AI agent systems [swamy2013verifying]. Based on this theory, one may also develop session types [gay1999types] for AI agents. The state monad has been utilized in building secure web applications [giffin2017hails] and microkernels [cock2008secure], and thus is a promising defense for the security of AI agents.
IV-B Sandbox
@@
A sandbox restricts the capabilities of the agent program. It enforces the limitation on the program’s access to both local and remote resources as shown in Figure 8. This section describes the application of classic access control provided by sandboxes on agent programs.
IV-B1 Access to local resources
The sandbox restricts the agent’s consumption of local resources such as CPU, memory, and storage. It also limits the agent’s access to a sub-file system. Together with session management, it further isolates the sub-file systems between sessions. To demonstrate the necessity of this approach, we designed BashAgent to interact with the operating system with bash as its tool, which uses gpt-3.5-turbo to understand user instructions and generate actions. BashAgent has two variants BashAgentf granted with full accessibility and BashAgentc constrained in a docker container. Based on AgentBench [liu2024agentbench], we collect and design 95 tasks related to system security to check the harmfulness of unconstrained AI agents. We categorize the tasks into confidentiality, integrity, and availability, and check if the LLM would accept the prompts with malicious intent and generate the attacking actions. We show the results of running BashAgentf in Table I. We found that BashAgentf accepts the majority $\nicefrac{{90}}{{95}}$ of malicious intents and generates the attacking instructions, and $\nicefrac{{76}}{{90}}$ generated attacking commands could be executed successfully in an unprotected environment, making the host system extremely vulnerable in all three security aspects. However, once we apply appropriate sandbox configurations, BashAgentc successfully defended against all the LLM-generated attacks. The LLM gpt-3.5-turbo was aligned with human values [ouyang2022training] but still struggles to reject malicious intent in the AI agent use case. Therefore, alignment training will not be enough to secure AI agents, and adding limitations on access to local resources is necessary for complete security.
TABLE I: Unconstrained AI agents will execute dangerous actions generated by the LLM. #Task is the number of tasks we gathered in this category. #Gen is the number of tasks accepted by the LLM and generates attacking actions. #Exec is the number of LLM-generated commends that are executed successfully and compromise the vulnerabilities.
@@ IV-B1 Access to local resources
IV-B2 Access to remote resources
Sandbox environment implements controlled access through mechanisms like whitelists, blacklists, and rate limiting in addition to fundamental interaction isolation. This framework allows resource providers to control the extent of access granted to agent programs selectively, ranging from full permission to complete prohibition or limitations to specific subsets of resources. Consequently, our method enhances security by effectively mitigating unwanted access from AI agents and potential threats posed by adversarial inputs to the agent.
IV-C Protecting Models for AI Agents
@@
IV-C1 Sessionless models for AI agents
@@
If the AI agent has no notion of sessions, then the agent must not fine-tune its LLM on private data or it must filter out private or malicious data from the query to the model.
@@
The first step is to identify this data. By employing meticulous prompt engineering, developers can enable the AI agent to interactively request sensitive data in a step-by-step manner, leaving markers on the data for further processing. The next step is to whitewash them into non-sensitive data. For example, by replacing US social security numbers (SSN) with nine random digits. This leaks no information about the specific SSN but still allows the model to learn from the context around the SSN. AI agent applications require this harmless version of data to be manipulable. For example, processing the last four digits of the credit card number as in web shopping [yao2022webshop]. In this case, the encryption transformation needs to be structure-preserving and information-preserving to text slicing. One solution for this is format-preserving encryption [bellare2010ffx].
Definition IV.1 (FPETS)
@@
A Format-Preserving Encryption for Text Slicing is an encryption scheme $E$ such that for all possible private messages $m$ and its indices $i,j$, $E(m[i\ldots j])=E(m)[i\ldots j],i\leq j$.
FPETS allows language models to read and manipulate private data as ciphertext instead of plaintext, therefore preventing privacy leaks. However, whether encrypting data in the input prompt harms the usability of the AI agent or not is unknown. To verify this defense method, we design an evaluation framework that prompts the LLM to operate on encrypted data. Each task in our evaluation framework is a roundtrip, where each AI agent is given a pair of encryption and decryption functions. When given a natural language prompt, the AI agents will first encrypt the data, and then pass the ciphertext to their LLM for manipulations such as text slicing. We then ask the agent to return the slice of information we want. The agent responds with the decrypted output for us to validate against the original slice of plaintext. We measure the success rate of this evaluation by $Succ=\nicefrac{{N^{\prime}}}{{N}}$ where $N$ is the total number of tasks and $N^{\prime}$ is the number of tasks where the agent completed a round trip with no error.
@@
As a proof of concept, we first tested encoded strings before encrypted strings. We generate random strings that include digits and both upper case and lower case letters, and encode them with a simple substitution cipher denoted by $E_{1}$, which extends the “rotate-by-13” cipher to operate on the character set mentioned above. Since $E_{1}$’s substitution on the characters is one-to-one, $E_{1}$ is FPETS. Let $D_{1}$ denote the decryption scheme corresponding to $E_{1}$. For confidential data $x$, this evaluation process can be formulated as $x=D_{1}(agent(E_{1}(x)))$.
@@
For comparison, we also report the success rate of the agent performing the same tasks with the plaintext in Table II. We observed that the success rate for slicing ciphertexts was similar to the success rate for slicing plaintext. Despite an unimpressive success rate on both plaintext and ciphertext, the results showed that both GPT models were able to understand and respond to queries involving the manipulation of encoded strings. Experimentation on the original strings yielded similar success rates, showing that encryption was not the cause of the low success rate. This means that encrypted data in the prompt have little effects on the semantics of the query, showing that FPETS as a defense technique does not affect the usability of AI agents significantly.
@@
TABLE II: Results for AI agent with encrypted data. Each agent is evaluated on 100 randomly-generated tasks. “SuccCiph” is the success rate of agent completing the tasks with encrypted data. “SuccPlain” is the success rate of the agent completing the same tasks without encryption.
@@
FPETS gpt-3.5-turbo 49.0% 47.0%
FPETS gpt-4-turbo 55.0% 57.0%
FHE gpt-3.5-turbo 85.0% 99.0%
FHE gpt-4-turbo 89.0% 94.0%
Text slicing is not the only task that an AI agent needs to complete on sensitive data. Another frequent use-case of AI agents is to perform calculations on sensitive data, which is common in financial and medical domains [naehrig2011can]. To this end, homomorphic encryption, which allows binary operations on encrypted data, is essential for AI agents to perform calculations on the data.
Definition IV.2 (FHE)
Let $\star$ be a binary operator. A homomorphic encryption scheme $\varphi:A\rightarrow B$ is a map from set of messages $A$ to $B$ such that for all $a,b\in A$, $\varphi(a\star b)=\varphi(a)\star\varphi(b)$. $\varphi$ is considered a fully homomorphic encryption scheme if it allows arbitrary function $\star$ to be applied to the data an unlimited number of times [acar2018survey].
Figure 9: Sessionless AI agents with encryption. Tools in this case need to be support a encryption scheme, like slicing for FPETS and addition or multiplication for FHE.
[…]
We report the evaluation results for FHE agents in Table II. Our evaluation results on addition and multiplication suggest that this defense is effective for AI agents requiring calculations on sensitive data supported by these operations. Thus, FHE is a solution for maintaining privacy during operations on sensitive data. Overall, our encryption defense does not substantially compromise the usability of AI agents and highlights a potential direction for future research on privacy-preserving AI agents.
IV-C2 Session-aware models for AI agents
An alternative to sessionless defenses is to make session-aware AI models. Towards this direction, OpenAI recently introduced Temporary Chat  [^2] https://help.openai.com/en/articles/8914046-temporary-chat-faq, where they promised not to use the chat history to improve their models. However, not improving the model on agent tasks would limit agent intelligence and user experience. To build powerful agent programs to handle diverse tasks, learning actions are essential.
One approach to privacy-preserving AI agents with personalization is fine-tuning each user’s LLM on their own chat history, isolating model updates per user as shown in Figure 10. However, this is costly and limited by available data. Alternatives like in-context learning [brown2020language] and retrieval-augmented generation [lewis2020retrieval] enhance responses by embedding past contexts in prompts, but are constrained by the length of model’s context window. A more promising method is prompt tuning [lester2021power], which freezes the foundational model and adds a few user-specific learnable parameters $\theta_{P}$ only to remember chat history. This technique avoids sharing data with the foundation model provider, directly addressing privacy concerns.
V Related work
Recent advancements in LLMs have had a significant impact in the development of AI agent, particularly in their ability to reason based on natural language prompts to observe and interact with their environments dynamically [wei2022chain, yao2024tree]. This shift from reinforcement learning to LLM agents has ushered in a new wave of AI agent development, where the emphasis is on enabling agents to perform actions based on natural language commands. ReAct [yao2023react] introduced chain-of-thought prompting [wei2022chain] to guide pre-trained LLMs to follow instructions in the agent setting. This approach has since been applied to computer tasks [kim2024language] and other real-world tasks [yao2022webshop, wang2023describe, gu2023dont, park2023generative]. To evaluate the performance of the agents, several benchmarks [zhou2024webarena, liu2024agentbench] have been proposed. These benchmarks measure the correctness of an agent’s actions without considering the potential vulnerabilities that agent actions can cause to the environment.
The threats to LLMs and AI agents are different [deng2024aiagentsthreatsurvey]. For LLMs, the concerns primarily address model alignment with human values, including ethics, offensive language, and politics [yu2024llm]. Conversely, AI agents, which use LLMs to generate actions and access tools, pose threats to real computing systems, applications, and resources, compromising their confidentiality, integrity, and availability.
VI Conclusion
With the aid of tool-augmented LLMs, AI agents are being recognized as a promising direction toward artificial assistants. Considerable research has focused on enhancing the accuracy of AI agent actions through advanced reasoning, planning, and learning. However, despite high performance in controlled evaluation settings, the potential side effects and dangers posed by these methods have not been thoroughly examined. In this paper, we present a systematic analysis of the security issues in current AI agent development and propose practical and feasible defense strategies. We discuss the potential vulnerabilities of AI agents both theoretically and in realistic scenarios with security-centric examples, and propose multiple defense techniques for each identified vulnerability. We highlight the future research directions and best practices for developing secure agent programs, and believe our work could boost the advancement of secure and trustworthy AI agents. Our code and data are publicly available  [^3] https://github.com/SecurityLab-UCD/ai-agent-security.
Acknowlegement
This work is partially supported by UC Noyce Initiative.
[…]
Appendix A Appendix
A-A Additional related work on system and software security
A-A1 Information security
However, these protection systems are compromised in the modern LLM-based agent systems, as AI agents’ behaviors are vastly different from regular user behaviors. For API-based LLMs like OpenAI’s GPT models, user contents, including input prompts and file uploads, are collected to improve services and develop new models. The models improved on user contents are vulnerable to training data leaks under specially designed adversarial attacks [nasr2023scalable]. Specialized fine-tuned agents also face this vulnerability, as the agent system may be shared by multiple users.
Encryption schemes can also keeping sensitive information secured [buchmann2004introduction, bishop2004introduction]. An encryption scheme is a 5-tuple ($P,C,K,\mathcal{E},\mathcal{D}$), where $P$ is the set of plaintext, $C$ is the set of ciphertext, $K$ is the set of keys, $\mathcal{E}$ is the set of encryption functions where for each $E_{i}\in\mathcal{E},E_{i}:P\times K\rightarrow C$, and $\mathcal{D}$ is the set of decryption functions, where for each $D_{i}\in\mathcal{D},D_{i}:C\times K\rightarrow P$. Encryption schemes allow people to freely exchange encrypted messages (ciphertexts) without revealing any private information (plaintexts) to unauthorized third parties who are not granted a key. Classical ciphers include transposition ciphers and substitution ciphers. These ciphers avert the aforementioned privacy leak vulnerability, yet limit AI agents’ ability to understand, process, or manipulate the data based on the special needs of the task. Homomorphic encryption [rivest1978data] is a family of encryption schemes that allow operations to be done on encrypted data without decrypting it first [acar2018survey].
A-A2 System security
The integrity of data on stored computing systems should also be secured for their accuracy. For this purpose, multiple policies with different focuses have been proposed [biba1977integrity, clark1987comparison]. Furthermore, data and resources should be available as they are needed. Beyond policy, isolation via virtualization is another common technique for access control. Sandboxes are environments where actions of a process are restricted by the policies [bishop2004introduction]. Sandboxes limit the access of the process on the system and therefore its consumption of computational resources and data [gong1989secure, fraser1998ensuring]. Over the years, different levels of sandbox and virtualization techniques have been created, including virtual machines, emulators, and containers [goldberg1974survey, qemu]. As a well-established method to limit computing resources and information accessibility in computer security [chiueh2005survey], virtualization protects the integrity of data and the availability of systems.
A-A3 Network security
A-B Additional Experiments
A-B1 FPETS on SSN
To demonstrate that Definition IV.1 works, we implement an SSN agent that manipulates nine-digit SSNs. In our design, all processes involving the plaintext of the SSNs are done outside of the underlying LLM. This ensures that the LLM is never exposed to raw sensitive information. We first provide the agent with an array of four secret keys and map each secret key to its respective randomly-generated SSN. The agent uses this information to encrypt each SSN and store the ciphertexts in an array. We then prompt the agent to return certain groups of an SSN, such as the first three or last four digits. Throughout its reasoning process, the underlying LLM of the agent has no access to the original plaintext of the SSNs. The LLM could at most call a tool to retrieve an SSN’s ciphertext for reasoning based on a fictitious user ID for indexing the array. After the LLM responds with the ciphertext representation of the slice we asked for, the agent replaces the ciphertext within the response with its decrypted value outside of the LLM before returning it to the user. We verified the results by comparing the agent’s response to the actual slice of plaintext we expected. The experimental results are shown in Table III.
TABLE III: Results for AI agent with encrypted SSN. Each agent is evaluated on 100 randomly-generated tasks. “SuccCiph” is the success rate of agent completing the tasks with encrypted data. “SuccPlain” is the success rate of the agent completing the same tasks without encrypting the data.
Agent Model SuccCiph SuccPlain